QUESTION 3 Let A = [ 4 0 1 −2 1 0 −2 0 1 ](its 3by3 matrix).

3.1. Find the eigenvalues of A.

3.2. For each eigenvalue of A find the corresponding eigenvectors.


3.3. Is A diagonalizable? Explain. If it is diagonalizable, then find P such that 𝐷 = 𝑃 −1𝐴𝑃 is a diagonal matrix. Do not compute 𝑃 −1(show all step!) . 


1
Expert's answer
2022-04-15T11:28:28-0400

"A = \begin{bmatrix} 4 & 0 & 1 \\ −2 & 1 & 0 \\ −2 & 0 & 1 \end{bmatrix}"

3.1. The characteristic polynomial of this matrix is

"P_{A}(\lambda)=\det (A-\lambda I) =\begin{vmatrix} 4-\lambda & 0 & 1 \\ −2 & 1-\lambda & 0 \\ −2 & 0 & 1-\lambda \end{vmatrix}="

"(1-\lambda)\begin{vmatrix} 4-\lambda & 1 \\−2 & 1-\lambda \end{vmatrix}=(1-\lambda)(\lambda^2-5\lambda+6)="

"=(1-\lambda)(\lambda-2)(\lambda-3)"

The eigenvalues of "A" are the roots of the polynomial "P_A(\lambda)":

"\lambda_1=1", "\lambda_2=2", "\lambda_3=3".


3.2. The eigenvector "X_1" corresponding to the eigenvalue "\lambda_1=1" is the solution to the equation "(A-\lambda_1I)X=0".

"\begin{bmatrix} 3 & 0 & 1 \\ −2 & 0 & 0 \\ −2 & 0 & 0 \end{bmatrix}\begin{bmatrix} x_1 \\ x_2 \\ x_3 \end{bmatrix}=\begin{bmatrix} 3x_1+x_3 \\-2 x_1 \\ -2 x_1 \end{bmatrix}=\begin{bmatrix} 0 \\0 \\0 \end{bmatrix}"

which implies "x_1=0", "x_3=0", and hence, "X_1=\begin{bmatrix} 0 \\1 \\0 \end{bmatrix}".

The eigenvector "X_2" corresponding to the eigenvalue "\lambda_2=2" is the solution to the equation "(A-\lambda_2I)X=0".

"\begin{bmatrix} 2 & 0 & 1 \\ −2 & -1 & 0 \\ −2 & 0 & -1 \end{bmatrix}\begin{bmatrix} x_1 \\ x_2 \\ x_3 \end{bmatrix}=\begin{bmatrix} 2x_1+x_3 \\-2 x_1-x_2 \\ -2x_1-x_3 \end{bmatrix}=\begin{bmatrix} 0 \\0 \\0 \end{bmatrix}"

which implies "x_1=t", "x_2=-2t", "x_3=-2t", and hence, "X_2=\begin{bmatrix} 1 \\-2 \\-2 \end{bmatrix}".

The eigenvector "X_3" corresponding to the eigenvalue "\lambda_3=3" is the solution to the equation "(A-\lambda_3I)X=0".

"\begin{bmatrix} 1 & 0 & 1 \\ −2 & -2 & 0 \\ −2 & 0 & -2 \end{bmatrix}\begin{bmatrix} x_1 \\ x_2 \\ x_3 \end{bmatrix}=\begin{bmatrix} x_1+x_3 \\-2 x_1-2x_2 \\ -2x_1-2x_3 \end{bmatrix}=\begin{bmatrix} 0 \\0 \\0 \end{bmatrix}"

which implies "x_1=t", "x_2=-t", "x_3=-t", and hence, "X_3=\begin{bmatrix} 1 \\-1 \\-1 \end{bmatrix}".


3.3. Consider the matrix "P=(X_1 | X_2|X_3)" (i.e. "X_1, X_2 ,X_3" are the columns of this matrix).

Then

"AP=(AX_1|AX_2|AX_3)=(\lambda_1X_1|\lambda_2X_2|\lambda_3X_3)=PD", where

"D=\begin{bmatrix} \lambda_1 & 0 & 0 \\ 0 & \lambda_2 & 0 \\ 0 & 0 & \lambda_3 \end{bmatrix}=\begin{bmatrix} 1 & 0 & 0 \\ 0 & 2 & 0 \\ 0 & 0 & 3 \end{bmatrix}" is a diagonal matrix.

Therefore, "D=P^{-1}AP", which means that the matrix "A" is diagonalizable.


Answer. The eigenvalues of "A" are "\lambda_1=1", "\lambda_2=2", "\lambda_3=3".

The eigenvalues of "A" are "X_1=\begin{bmatrix} 0 \\1 \\0 \end{bmatrix}" , "X_2=\begin{bmatrix} 1 \\-2 \\-2 \end{bmatrix}", "X_3=\begin{bmatrix} 1 \\-1 \\-1 \end{bmatrix}".

The matrix "P" equals to "(X_1 | X_2|X_3)=\begin{bmatrix} 0 & 1 & 1 \\ 1 & -2 & -1 \\ 0 & -2 & -1 \end{bmatrix}".